OpenStudy (anonymous):

What is the missing number? Five-sixths is equal to how many twenty-fourths? 4 5 8 twenty

OpenStudy (anonymous):

Set two fractions equal and cross multiply.

OpenStudy (anonymous):

(5/6) = (x/24) 6x = 5(24) 6x = 120 divided by 6 on both sides. x = 20

OpenStudy (anonymous):

Another way, although slower way of checking is by going through all the answers: (5/6) = 0.833333333 (4/24) = 0.16666666 (5/24)=0.208333333 (8/24)=0.333333333 (20/24)=0.83333333=(5/6)

OpenStudy (anonymous):

but the answer should be 25/4 right or no??

OpenStudy (anonymous):

or sorry no lmao im on a different question wth